OpenStudy (smileyxl3):

If g(t)=(3t-5)^2+4t-1 find expressions for the following

OpenStudy (smileyxl3):

a. g(a) b. g(a+2)

OpenStudy (smileyxl3):

@satellite73

OpenStudy (smileyxl3):

@hartnn

OpenStudy (smileyxl3):

@AravindG

OpenStudy (lucaz):

I guess you just have to substitute t by a and (a+2)

OpenStudy (lucaz):

like if they were values of t

OpenStudy (smileyxl3):

@lucaz I don't know... well at least you answered.

OpenStudy (lucaz):

g(t) you pick a number, 3, g(3)=(3(3)-5)^2+4(3)-1 if you pick a variable just replace the variable, or expression

OpenStudy (lucaz):

g(a)=(3(a)-5)^2+4(a)-1 or g(a+2)=(3(a+2)-5)^2+4(a+2)-1